# MultiCrafter: High-Fidelity Multi-Subject Generation via Disentangled Attention and Identity-Aware Preference Alignment

### v1

@@ Appendix A More Details for Identity-Preserving Preference Optimization
To advance research on high-fidelity multi-subject generation, we constructed a benchmark dataset by collecting images from publicly available sources and extracting the corresponding facial regions as reference images. The dataset comprises 80 celebrities and 80 non-celebrities, covering diverse attributes in terms of gender, age, and ethnicity (male/female; young/elderly; Caucasian, Black, and Asian). We used this face collection as a reference pool and paired faces within it. For each pair, we employed Qwen2.5-VL to generate distinctive natural-language prompts to provide diverse textual descriptions. Representative samples are shown in Fig.˜6.
@@
First, the scale and quality of the training data are the primary limiting factors. Currently, high-quality, publicly available datasets for multi-subject driven generation remain scarce (Wu et al., 2025c; Chen et al., 2025a). Although we have designed a complete automated data processing pipeline to extract training samples from videos, our dataset is still limited in scale and diversity due to the quantity and quality of open-source video data.
@@
Second, the effectiveness of our method has been primarily validated in two-subject scenarios.Since our multi-person dataset and the public MUSAR dataset (Guo et al., 2025) mainly contain two subjects, the experiments in this paper were centered around this setting. The model’s generation capability when handling three or more subjects has not been fully verified. It is worth noting that our framework was designed with scalability in mind; both the attention regularization mechanism and the Multi-ID Alignment Reward (based on the Hungarian algorithm) in the reinforcement learning framework can be directly extended to scenarios with more subjects.
@@
For future work, we plan to explore improvements from both data and model perspectives. On one hand, we will attempt to construct larger, higher-quality datasets containing a more diverse number of subjects by combining synthetic data with image editing Wu et al. (2025a) techniques. On the other hand, we will train and evaluate the model in scenarios with more subjects to further enhance the generalization and robustness of MultiCrafter, enabling it to handle more complex personalized image generation.
Appendix F The Use of Large Language Models.
In this paper, we only use the large language model to help polish our text. The large language model has no role in the research conception.
Appendix G More Results of Multi-Human Personalization.
@@
To further demonstrate our method’s performance in multi-human personalization, we present qualitative comparisons in Fig.˜7 and Fig.˜8. The results show that our model effectively preserves the identity of each subject and avoids the "attribute leakage" common in other methods. This outcome validates the efficacy of our Identity-Disentangled Attention Regularization (IDAR). While some baselines produce more stylized outputs that may yield higher HPSv2 scores, this is often at the expense of subject fidelity. Our method prioritizes photorealism and faithful subject appearance consistency, which leads to more reliable results in multi-subject customization.
Figure 7: More Visualization of our method in Multi-Human Personalization.
Figure 8: More Visualization of our method in Multi-Human Personalization.
Appendix H More Results of Multi-Object Personalization.
@@
To evaluate the generalization of our framework, we showcase multi-object customization comparisons in Fig.˜9. Our method demonstrates high object fidelity, accurately preserving core visual attributes such as a toy’s texture or a glass’s geometry. In contrast, competing approaches often introduce artifacts like deformation and detail loss. This highlights our model’s strength in precise subject representation rather than hyper-stylization, a crucial capability for practical applications that require accuracy.
Figure 9: Visualization of our method in Multi-Object Personalization.
Appendix I More Results of Single-Subject Personalization.
@@
Effective multi-subject generation builds on strong single-subject performance. We validate this capability in Fig.˜10 and Fig.˜11, showing six diverse samples for each of four individuals and six frontal single-subject comparisons against SOTA models. Our method consistently preserves identity across varying styles, poses, and scenes, and even improves fidelity to the reference over baselines. These results confirm that the proposed framework not only enables reliable multi-subject generation but also enhances single-subject identity fidelity.
Figure 10: Visualization of our method in Single-Subject Personalization.

### v2

@@ Appendix D Impact Analysis of Dataset Quality
To advance research on high-fidelity multi-subject generation, we constructed a benchmark dataset by collecting images from publicly available sources and extracting the corresponding facial regions as reference images. The dataset comprises 80 celebrities and 80 non-celebrities, covering diverse attributes in terms of gender, age, and ethnicity (male/female; young/elderly; Caucasian, Black, and Asian). We used this face collection as a reference pool and paired faces within it. For each pair, we employed Qwen2.5-VL to generate distinctive natural-language prompts to provide diverse textual descriptions. Representative samples are shown in Fig.˜3.
@@
First, the scale and quality of the training data are the primary limiting factors. We discussed in detail the limitations imposed by the quality of our data in  Appendix˜D. Currently, high-quality, publicly available datasets for multi-subject driven generation remain scarce [59, 3]. Although we have designed a complete automated data processing pipeline to extract training samples from videos, our dataset is still limited in scale and diversity due to the quantity and quality of open-source video data.
@@
Second, the effectiveness of our method has been primarily validated in two-subject scenarios. Since our multi-person dataset and the public MUSAR dataset [17] mainly contain two subjects, the experiments in this paper were centered around this setting. Although we have supplemented some results from a small number of customized generation experiments with 3 person in  Appendix˜I, the model’s generation capabilities when dealing with more objects have not yet been fully validated. It is worth noting that our framework was designed with scalability in mind; both the attention regularization mechanism and the Multi-ID Alignment Reward (based on the Hungarian algorithm) in the reinforcement learning framework can be directly extended to scenarios with more subjects.
@@
For future work, we plan to explore improvements from both data and model perspectives. On one hand, we will attempt to construct larger, higher-quality datasets containing a more diverse number of subjects by combining synthetic data with image editing [57] techniques. On the other hand, we will train and evaluate the model in scenarios with more subjects to further enhance the generalization and robustness of MultiCrafter, enabling it to handle more complex personalized image generation.
Appendix G User Study
Figure 4: User Study. Our method achieved the best results in both subject fidelity and overall quality, demonstrating its effectiveness.
To complement our quantitative analysis and assess human subjective preference, we conducted a user study comparing MultiCrafter against five state-of-the-art baselines: UNO [59], OmniGen [62], OmniGen2 [58], DreamO [43], and XVerse [3]. We evaluated the generated results across five key dimensions: text alignment, subject fidelity, aesthetics, realism, and overall preference. As illustrated in LABEL:fig:user_study, our method demonstrates a commanding lead in subject fidelity, significantly outperforming the second-best method. This result strongly validates the effectiveness of our Identity-Disentangled Attention Regularization in resolving attribute leakage and preserving intricate identity details. Furthermore, MultiCrafter achieves the highest scores in realism and overall quality, while maintaining a top-tier performance in text alignment and competitive aesthetics. These findings confirm that our decoupled training framework successfully resolves the trade-off between fidelity and preference, producing images that are not only faithful to the user-provided subjects but also aesthetically pleasing and semantically accurate.
Appendix H More Discussion about MoE-LoRA.
To further validate the effectiveness of our MoE-LoRA architecture in expanding model capacity and handling diverse spatial layouts, we analyzed the expert routing behavior under distinct spatial instructions. If the MoE mechanism truly functions as intended, the routing network should dynamically allocate different experts to handle different spatial layouts. We designed a controlled experiment using two distinct spatial prompts: a ‘side-by-side’ (left-right) layout and a ‘top-down’ (up-down) layout. To isolate the impact of spatial configuration, we kept all other variables constant, including the subject identities, background descriptions, and random seeds. We conducted inference across 4 different seeds and calculated the average change ratio of the expert routing distribution within the Double Blocks. The change ratio for an expert $e$ at layer $l$ is defined as $|P_{A}^{(l,e)}-P_{B}^{(l,e)}|/(P_{A}^{(l,e)}+\epsilon)$, where $P_{A}$ and $P_{B}$ represent the routing probabilities under the two different layouts.
Figure 5: Layer-wise Change Ratio in Expert Routing Distribution. The heatmap illustrates the difference in expert activation between "Left-Right" and "Top-Down" layouts.
The results are visualized in Fig.˜5. As shown in the heatmap, significant variations in expert activation (indicated by deep blue regions, where the change ratio exceeds $40\%-50\%$) are observed across many layers. This phenomenon proves that the routing network effectively perceives the change in spatial prompts. The distinct activation patterns confirm that the model utilizes different experts to construct different layouts. This experiment strongly supports our claim that MoE-LoRA successfully decouples the learning of diverse spatial scenarios. By assigning specialized experts to different spatial layouts, the framework avoids the "attribute averaging" issue inherent in single-LoRA approaches, thereby achieving high fidelity across complex and varied compositions. In addition, we have added more visualizations of the results for multi-human and multi-subject generation. As shown in Fig.˜6, after adding MoE-LoRA, many scenes that could not be generated before adding MoE-LoRA can be generated accurately.
Figure 6: More Visualization of the effectiveness of MoE-LoRA.
Appendix I More Results of Multi-Human Generation.
@@ Appendix G User Study
To further demonstrate our method’s performance in multi-human personalization, we present qualitative comparisons in Fig.˜7 and Fig.˜8. The results show that our model effectively preserves the identity of each subject and avoids the "attribute leakage" common in other methods. This outcome validates the efficacy of our Identity-Disentangled Attention Regularization (IDAR). While some baselines produce more stylized outputs that may yield higher HPSv2 scores, this is often at the expense of subject fidelity. Our method prioritizes photorealism and faithful subject appearance consistency, which leads to more reliable results in multi-subject customization. Furthermore, we extended our method to scenarios involving more subjects to validate the effectiveness of our framework. We curated 10,000 samples of three-person customized generation data from video data to train with our method. As shown in Fig.˜9, our method continues to significantly improve subject fidelity compared to existing methods, while maintaining text alignment and image aesthetics, effectively aligning with multi-dimensional human preferences. This demonstrates the effectiveness of our framework.
Figure 7: More Visualization of our method in Multi-human Generation.
Figure 8: More Visualization of our method in Multi-human Generation.
Figure 9: More Visualization of our method in three-human Generation.
Appendix J More Results of Multi-Object Generation.
@@
To evaluate the generalization of our framework, we showcase multi-object customization comparisons in Fig.˜10. Our method demonstrates high object fidelity, accurately preserving core visual attributes such as a toy’s texture or a glass’s geometry. In contrast, competing approaches often introduce artifacts like deformation and detail loss. This highlights our model’s strength in precise subject representation rather than hyper-stylization, a crucial capability for practical applications that require accuracy.
Figure 10: Visualization of our method in Multi-object Generation.
Appendix K More Results of Single-Subject Generation.
@@
Effective multi-subject generation builds on strong single-subject performance. We validate this capability in Fig.˜11 and Fig.˜12, showing six diverse samples for each of four individuals and six frontal single-subject comparisons against SOTA models. Our method consistently preserves identity across varying styles, poses, and scenes, and even improves fidelity to the reference over baselines. These results confirm that the proposed framework not only enables reliable multi-subject generation but also enhances single-subject identity fidelity.
Figure 11: Visualization of our method in Single-subject generation.